OpenStudy (tia_sweetheart17):

X^2 divided by x over b

OpenStudy (agreene):

x^2 is the same as x*x, so \[\frac{x*x}{x}\] x cancels x giving us \[\frac{x}{b}\] And, because we know a thing or two about fractions, we can also say: \[\frac{x}{b}:b\ne0\]